OpenStudy (anonymous):

becca's brownie recipe calls for 3 3/4 cups of flour. if becca wants to make a batch of brownies that is 3/5 amount of the standard recipe, how much flour will she need?

OpenStudy (anonymous):

Just take (3/5) * (15/4) = 9/4 cups (= 2 1/4)

OpenStudy (anonymous):

and 15/4 of course equals 3 3/4